OpenStudy (anonymous):

Help! Will give metal! Find the length of the base of a square pyramid if the volume is 128 cubic inches and has a height of 6 inches. Use 3.14 for pi and round your answer to the nearest hundredth. A. 4 inches B. 8 inches C. 16 inches D. 24 inches I know you're supposed to divide 128 by 6 but then what?

OpenStudy (fibonaccichick666):

first and foremost and always, draw a picture

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

there

OpenStudy (anonymous):

you may have to click on it it's a bit small

OpenStudy (fibonaccichick666):

alright, I'll let that pic slide. next what is the formula for volume of a square pyramid?

OpenStudy (anonymous):

haha I had to do it on paint real quick sorry. The formula for the volume of a square pyramid is V=1/3(b)(h)

OpenStudy (anonymous):

she left...

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (anonymous):

She came :))

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

yup!

OpenStudy (fibonaccichick666):

alright so now, you have the height, the volume, and you are looking for b(which I THINK YOU MAY HAVE MISWRITTEN) BECAUSE WHAT DOES THAT b stand for?

OpenStudy (fibonaccichick666):

sorry caps lock

OpenStudy (anonymous):

it stands for the base of the pyramid

OpenStudy (fibonaccichick666):

it does not, actually, use this awesome song to help you find out what it does stand for :D https://www.youtube.com/watch?v=xHQHuWZBZlY

OpenStudy (anonymous):

k watching it

OpenStudy (fibonaccichick666):

kk

OpenStudy (fibonaccichick666):

it has V=Bh, but same gist

OpenStudy (anonymous):

oh ok it stands for area of the base

OpenStudy (fibonaccichick666):

yup yup awesome song right???? I love it!

OpenStudy (anonymous):

haha ya

OpenStudy (fibonaccichick666):

so what does the area of the base = in this case?

OpenStudy (anonymous):

hmm I'm not sure the problem has 3.14 in it but I'm not sure what that's used for

OpenStudy (anonymous):

either

OpenStudy (fibonaccichick666):

well, what is the area of a square?

OpenStudy (fibonaccichick666):

and that's extraneous info

OpenStudy (anonymous):

Sorry google chrome crashes on my computer sometimes so that's why it took me so long to reply. The area of a square is the base * the height correct?

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (fibonaccichick666):

no

OpenStudy (fibonaccichick666):

draw a square

OpenStudy (fibonaccichick666):

lbel the sides some variable, then tell me the area

OpenStudy (anonymous):

s^2?

OpenStudy (fibonaccichick666):

alright so now, what is the area of the base of a square pyramid if the side lingtgh is b?

OpenStudy (fibonaccichick666):

length*

OpenStudy (fibonaccichick666):

then what is our new formula for volume?

OpenStudy (anonymous):

s^2 since the base of a square pyramid is a square? so would it be V=1/3(s^2)(h)

OpenStudy (fibonaccichick666):

yup yup

OpenStudy (fibonaccichick666):

so now plug in your values and solve for s

OpenStudy (anonymous):

128in^3=(s^2)(6).. divide both sides by 6 so now 21.33=(s^2) but then what? I don't know what's next do i subtract, add, divide? I'm really confused.

OpenStudy (fibonaccichick666):

how do you undo s times s aka s^2?

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

so i dived 21.33 by 2?

OpenStudy (anonymous):

divide*

OpenStudy (fibonaccichick666):

no, you use something called a square rtoot

OpenStudy (fibonaccichick666):

root*

OpenStudy (fibonaccichick666):

√(s^2)=s

OpenStudy (anonymous):

mmk

OpenStudy (anonymous):

21.33=s

OpenStudy (anonymous):

but that's no in any of the answers

OpenStudy (anonymous):

not*

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (anonymous):

am I supposed to add 3.14 now?

OpenStudy (anonymous):

yup

OpenStudy (fibonaccichick666):

no

OpenStudy (fibonaccichick666):

i told you you don't need pi

OpenStudy (anonymous):

I don't know math is very hard for me. I'm doing my best sorry

OpenStudy (fibonaccichick666):

you did the problem. The answers and question do not match up however

OpenStudy (fibonaccichick666):

OH WAIT, you forgot the 1/3

OpenStudy (anonymous):

xD i forgot about it too

OpenStudy (fibonaccichick666):

currently you have 128/6=1/3 *s^2

OpenStudy (anonymous):

divide by 1/3 is the next step right?

OpenStudy (fibonaccichick666):

yea

OpenStudy (anonymous):

ok I did that before but i wasn't sure since it equals 7.1 and that's not on the list so the closest answer to it is 8

OpenStudy (fibonaccichick666):

128/2=64 right..?

OpenStudy (anonymous):

yes

OpenStudy (fibonaccichick666):

you miscalculated check your work

OpenStudy (fibonaccichick666):

it's an even answer

OpenStudy (fibonaccichick666):

a whole number

OpenStudy (anonymous):

128/6= 21.3/3=7.1 what am I miscalculating

OpenStudy (fibonaccichick666):

that isn't dividing by 1/3 that's multiplying by it

OpenStudy (anonymous):

i get 64 but that not on the list

OpenStudy (anonymous):

or that's what my calculator says anyways

OpenStudy (fibonaccichick666):

64=?

OpenStudy (fibonaccichick666):

you are still solving for s...

OpenStudy (anonymous):

i know I divide 21.3 by 1/3 and i get 64

OpenStudy (anonymous):

k I'm super lost now nvm I give up thanks for trying to help me I'll still give you a medal though

OpenStudy (anonymous):

if anyone checks this Question out cause they're doing homework or what not just know the answer isn't 24.

OpenStudy (fibonaccichick666):

no, but stop you are so darn close

OpenStudy (fibonaccichick666):

you divided what by 1/3 why? what were you solving what were you getting by itself?

OpenStudy (anonymous):

i divided 21.33 by 1/3

OpenStudy (anonymous):

getting s by itself

OpenStudy (fibonaccichick666):

so what does 64 =? what was the other side what was left over there?

OpenStudy (anonymous):

64=s?

OpenStudy (fibonaccichick666):

remember we were solving 128/6=1/3 *s^2

OpenStudy (anonymous):

yes ok so 128/6/one third=s^2

OpenStudy (anonymous):

s^2 is still left

OpenStudy (fibonaccichick666):

yea

OpenStudy (fibonaccichick666):

so 64=s^2

OpenStudy (fibonaccichick666):

now s=?

OpenStudy (anonymous):

divided 64 by 2?

OpenStudy (anonymous):

well s^2

OpenStudy (fibonaccichick666):

no, remember how do you undo a square?

OpenStudy (anonymous):

sqrt

OpenStudy (fibonaccichick666):

yup

OpenStudy (anonymous):

find the sqrt of 64 or s^2

OpenStudy (fibonaccichick666):

both

OpenStudy (anonymous):

sqrt of 64 is 8